A total of $10,000 is invested in two funds paying 5% and 7% annual interest. The combined annual interest is $630. How much of the $10,000 is invested in each fund?
Answer by mananth(16946)

Investment I 5.00% per annum
Investment II 7.00% per annum

x + y 10000 ------------------------1
5.00% x + 7.00% y = $630.00
Multiply by 100
5 x + 7 y = $63,000.00 --------2

Multiply (1) by -5
we get

-5 x -5 y = -50000.00

Add this to (2)

0 x +2 y = $13,000.00

divide by 2

y =$6,500.00 investment at 7.00%
Balance $3,500.00 investment at 5.00%
CHECK
$3,500.00 $175.00
$6,500.00 $455.00
Total $630.00